# Auto Arena of LLMs: Automating LLM Evaluations with Agent Peer-battles and Committee Discussions

### v3

@@ 4.2 Results: Alignment with Human Preferences
Table 1: Correlation analysis with Chatbot Arena of evaluation benchmarks on 7 LLMs.
Spearman Correlation
MMLU 89.3%
OpenLLM 89.3%
MT-Bench 82.9%
MT-Bench-Hard 89.3%
LC-AlpacaEval 90.0%
Auto-Arena 96.4%
We regard Chatbot Arena scores as a trustworthy indicator of human preferences and general capabilities of LLMs. Table 1 shows the Spearman correlations with Arena scores achieved by various benchmarks. As all benchmarks are evaluated only in English, we use English-only Arena scores. We see that both static and model-based baselines result in a similar level of correlation no higher than 90%, with LC-AlpacaEval slightly surpassing others. Then, Auto-Arena can improve the correlation to 96.4%, outperforming SOTA by 6.4%. Notably, among all benchmarks, Auto-Arena is the only one that requires no human efforts, either on dataset compilation or judgment generation. The high alignment with human preferences could originate from the human-like design, which effectively mimics the human users’ voting processes.
Table 2: Spearman Correlation between Auto-Arena and Chatbot Arena (“Arena”) With and Without Peer Battles.
Auto-Arena to Arena
Without Peer Battles 50.0%
With Peer Battles 96.4%
@@
Auto Arena (Before discussion) 48%
Auto Arena (After discussion) 68%
@@
4.3 Ablation Studies on Peer Battles and Committee Discussions
To investigate the effect of "peer battles" on the evaluation quality, we conduct an ablation study and show the results in Table 3. As a baseline, we ask the committee to collectively evaluate the two candidates’ initial responses to the raised question. Without the peer battle component, the correlation with human preferences drops from 96.4% to 50.0%. This proves the importance of the peer battles, during which the performance gaps between candidates become more visible and robust to judges. Thus, peer battles can improve evaluation reliability and alignment with human preferences.
Figure 4: Cohen’s Kappa Agreement with Majority Vote After Committee Discussions.
The committee discussion component is designed to introduce diverse viewpoints and produce more consistent verdicts. As shown in Figure 4, the Cohen’s Kappa agreement [29] between individual judges and the final result (voted) is low, averaging 0.36, which indicates fair agreement. Specifically, compared to strong models, the judgments of weak models align less with the voted result, such as Llama-2 compared to GPT-4. This shows that general model capabilities could result in significant performance gaps when used as judges. After the Committee Discussion, Figure 4 shows an overall improvement in agreement, averaging 0.56, which indicates moderate agreement. In the discussion process, judges are exposed to more viewpoints, among which some may be convincing enough to result in a change in verdict. More analysis on inter-judge agreement is provided in Appendix A.3.
Table 3 shows the agreement probability among judges. Agreement probability is defined as the mean probability of two random judges agreeing with each other. After committee discussion, the agreement increases by 20%, matching the agreement level among human annotators on MT-Bench. This observation indicates that committee discussions can significantly improve the quality of judgments in terms of consistency and reliability, matching human-level performance.
5 Constructing a Leaderboard with Auto-Arena
Figure 5: Elo Score Changes of Adding Llama-3 to the Ranking of 7 Models.
Figure 6: Elo Scores of 17 Models by Auto-Arena on English.
@@ 5.1 Update New Models to Leaderboard
With Auto-Arena, we can obtain the rank for a list of models with their ELO scores to construct a leaderboard. As new LLMs come out frequently, we describe how to add new candidate models to the existing leaderboard with 10 more models that are released very recently. To add a new candidate, we ask it to debate with $\lceil log_{2}(n)\rceil$ opponents with similar ELO scores, where $n$ is the number of total participants after adding the new candidate. We initialize the first pairing by asking the new candidate to debate with the opponent with the most similar MMLU score. This addition mechanism is generalizable and maintains the computational costs of evaluating $n$ models below $nlog_{2}(n)$.
As an example, we add a new participant (Llama-3) to the existing 7-model ranking. It battles with $\lceil log_{2}(8)\rceil=3$ close opponents and Figure 6 shows the ELO score changes. Firstly, it is paired with Qwen-1.5 based on MMLU similarity and wins, which results in a very high ELO score, even above GPT-4. Then, it is paired with GPT-4, the most similar opponent. After losing, the score stabilizes at second place. Finally, it is paired with Claude-3-Haiku, the closest opponent besides GPT-4. It wins and secures the final ranking in second place. This process lets the new candidate battle with a reasonable fraction of close opponents and makes the final ranking stable without disrupting the other participants, whose score distribution remains similar before and after the addition.
Table 4: Correlation analysis with Chatbot Arena of evaluation benchmarks on 17 LLMs after extension. The 4 models not on Chatbot Arena are skipped.
@@
MMLU 90.1%
OpenLLM 85.7%
MT-Bench 89.3%
MT-Bench-Hard 86.7%
LC-AlpacaEval 90.3%
Auto-Arena 94.5%
@@
Using this scalable addition approach, we construct a comprehensive leaderboard by adding 10 models to the existing tournament. Besides the best-performing and newest models from each major model family in the top 20 list on Chatbot Arena (detailed in Appendix A.4), we add 4 popular Chinese LLMs, which are GLM [^5] https://open.bigmodel.cn/, SenseChat [^6] https://platform.sensenova.cn/home, Minimax [^7] https://platform.minimaxi.com/examination-center/text-experience-center, and Wenxin [^8] https://cloud.baidu.com/wenxin.html. Figure 6 shows the overall ELO scores by Auto-Arena on the 17 models. On the 13 models included in Chatbot Arena, we can recover rankings with 92.3% accuracy. Table 7 shows the Spearman correlations after expansion. Auto-Arena remains the method most aligned with human preferences by a margin of 4%. Therefore, Auto-Arena is generalizable and robust for maintaining a leaderboard on a large number of LLMs.
5.2 Easy Extension to Other Domains and Languages
As Auto-Arena of LLMs is fully automatic, it can be easily adapted to evaluate LLMs in non-mainstream domains or languages. As a case study, we conduct a Chinese tournament on 11 out of the 17 models that are claimed to have multi-lingual proficiency. The only adaptation effort required is translating the prompts into Chinese. Then, the examiner automatically generates questions in Chinese and the candidates battle in Chinese. Similarly, for adaptation to another domain, the only effort is to change the “domain” specification in the examiner’s prompts.
Figure 7 shows the ELO scores derived by Auto-Arena for Chinese evaluation. We notice that the ranking differs significantly from English results. For example, SenseChat-5, an LLM specifically trained for Chinese usage, is able to improve from 10th place (in English) to 2nd place (in Chinese). On the contrary, Llama-3-70b, which only uses 5% multilingual data during pretraining [30], drops from 4th place (in English) to 10th place (in Chinese). As Chinese evaluation benchmarks are limited, we compare to the Chinese-only leaderboard on Chatbot Arena, which constitute for 10.36% of all collected dialogues. On the 7 models that are also included in Chatbot Arena, Auto-Arena recovers their ELO scores with 92.86% correlation and restores rankings with 90.5% accuracy.
@@ 6 Investigation of LLM’s Behaviors in Competitive Peer Battles
The example in Figure 9 shows excerpts of a peer battle around question: “how many unique ways to arrange letters in ‘LETTER’.” Candidate A (powered by Yi-34B-Chat) gives a wrong answer as it miscounts occurrences for repeated letters and miscalculates factorials. The opponent B (powered by Claude-3-Haiku) quickly and precisely points out these two issues and skillfully raised a follow-up that targets A’s weaknesses: “how about the word ‘BANANA’?” Then, A still miscalculates factorials. Seeing these results, the judge then determines that B is the better assistant. In peer battles, we see that LLM candidates efficiently understand the rules of the competitive environment and can design targeted strategies to attack the opponent in order to win. The peer battle process then ensures that each raised question is tailored to the specific LLM, resulting in more specialized evaluations.
@@
Peer Battles Make the Performance Gaps Become Visible
In the example shown in Figure 10, given a math question on infinite series, both candidate A (Claude-3-Haiku) and candidate B (GPT-4-Turbo) provide correct answers in the first round. However, as the debate deepens, the performance gap becomes more visible: Candidate B is able to provide a more elaborate and helpful response. In the ablation study without peer-battles, the judges initially decided that it was a tie. However, after seeing the subsequent debates, they change to favoring assistant B. This example shows that the debate process indeed pushes the candidate LLM’s capabilities to the limit, testing deeper understandings and reasoning abilities. Also shown in the previous Table 3, the peer-battle design is indispensable for a robust and comprehensive evaluation.
@@ 7 Conclusions
In this paper, we innovatively design a completely automatic evaluation framework: Auto-Arena of LLMs. By using LLM agents to generate questions, employing LLM candidates in peer battles, and evaluating responses using LLM committee discussions, Auto-Arena produces less-contaminated, timely, and trustworthy evaluations. In the extensive experiments, Auto-Arena achieves the highest correlation with human preferences, despite requiring zero human efforts. It is easily adaptable to other domains and resources, promoting fairness of AI system evaluations. The peer battles also demonstrate several interesting LLM behaviors in competitive environments, including attacking and learning from the opponents. Moreover, there still exist some limitations to the current approach. For example, the distribution of question domains is artificially designed, which may deviate from real-life distributions. Currently, Auto-Arena focuses on 1-to-1 peer battles, which limits its usage in multi-player scenarios. Along with the paper, we release a website, demos, and an actively-maintained leaderboard. We hope this work could serve as a valuable tool for LLM evaluation.
Appendix A Appendix
A.1 Prompts Used
@@
A.1.1 Prompts to Examiner agent
@@
A.1.2 Prompts to Peer Battle Candidates
@@
This is the beginning prompt to the peer battle candidates. When possible, it is included as a system prompt. The action guide prompts are included in Table 6, where the actions are determined by the round and turn as illustrated in Figure 2.
@@ A.1.3 Prompts to Judges
This is the prompt to judges for discussion:
@@
A.2 Example Questions Generated
@@ 2. What is the integral of ($x^{2}+2x+2)/(x^{3}+3x^{2}+3x+1)dx$?
A.3 Inter-judge Agreement

### v4

@@ 3.2 Results: Alignment with Human Preferences
We regard Chatbot Arena scores as a trustworthy indicator of human preferences and general capabilities of LLMs. Table 2 shows the Spearman correlations with Chatbot Arena scores achieved by various benchmarks. As all benchmarks are evaluated only in English, we use English-only Chatbot Arena scores. We see that both static and model-based baselines result in a similar level of correlation that is below 90%, with Arena-Hard surpassing others at 85.71%. Then, Auto-Arena can improve the correlation to 91.67%, outperforming the SOTA by 5.96%. Notably, among all benchmarks, Auto-Arena is the only one that doesn’t require human efforts, neither on dataset compilation nor judgment generation. The high alignment with human preferences could originate from the human-like design, which effectively mimics the human users’ voting processes.
3.3 Ablation Studies on Peer Battles and Committee Discussions
Peer-battles:   We conduct an ablation study on whether peer-battles affect the evaluation quality and include the results in Table 2 (“w/o Peer Battles”). In this setup, we ask the committee to only evaluate the two candidates’ initial responses to the synthetic question, where the judge prompts stay the same. For this no-debate design, the question-answering process mimics that of MT-Bench or LC-AlpacaEval, but with an added committee discussion component. As a result, we observe that the correlation is slightly higher than LC-AlpacaEval and MT-Bench by a margin of 3.81%. Compared to the full Auto-Arena framework, however, the performance drops by 5.00%. This proves the effectiveness of the peer battles, during which the performance gaps between candidates become more visible and robust to judges. Thus, peer battles can improve alignment with human preferences.
Figure 3: Cohen’s Kappa agreement with majority vote results before (upper) and after (lower) committee discussions.
@@ 3.3 Ablation Studies on Peer Battles and Committee Discussions
Auto-Arena (Before discussion) 53%
Auto-Arena (After discussion) 64%
@@
Committee Discussions:   The committee discussion component is designed to introduce various points of view and produce more consistent decisions. As shown in Table 2, the correlation with human preferences drops from 91.67% to 88.33% without committee discussions, showing the effectiveness of the component in improving evaluation quality. As shown in Figure 3, before committee discussions, the Cohen’s Kappa agreement (McHugh, 2012) between individual judges and the final result (voted) is low, averaging 0.41. Specifically, compared to strong models, the judgments of weak models align less with the voted result, such as Yi compared to GPT-4. This shows that general model capabilities could result in significant performance gaps when used as judges. After the committee discussions, agreement increased to an average of 0.54, which indicates moderate agreement. In the discussion process, judges are exposed to more viewpoints, among which some may be convincing enough to result in a change in verdict. More analysis on the inter-judge agreement is provided in Appendix F, where we see that discussions could largely improve the agreements among individual judges as well. Table 3.3 shows the agreement probability among judges. Agreement probability is defined as the mean probability of two random judges agreeing with each other. After committee discussion, the agreement increases by 11%, matching the agreement level among human annotators on MT-Bench. This observation indicates that committee discussions can significantly improve the quality of judgments to match with human-level performance.
4 Constructing and Maintaining a Leaderboard with Auto-Arena
Figure 4: Changes in Elo scores of adding Llama-3 to the ranking of 9 models.
Figure 5: Elo scores of 15 models by Auto-Arena on English.
@@ 4.1 Update New Models to Leaderboard
With Auto-Arena, we can obtain the rank for a list of models with their Elo scores to construct a leaderboard. As new LLMs are released frequently, we describe how to add new candidate models to the existing leaderboard with 6 more models which are released very recently, as previously listed in Section 3.1. To add a new candidate, we ask it to debate with $\lceil log_{2}(n)\rceil$ opponents with similar Elo scores, where $n$ is the number of total participants after adding the new candidate. For the first pairing, as we do not have Elo indicators, we initialize by asking the new candidate to debate with the opponent with the most similar MMLU score. This addition mechanism is generalizable and maintains the computational costs of evaluating $n$ models below $nlog_{2}(n)$.
Table 4: Correlation analysis with Chatbot Arena of evaluation benchmarks on 15 LLMs after extension.
@@
OpenLLM 32.50%
GPQA 62.86%
MMLU 46.20%
LC-AlpacaEval 76.32%
MT-Bench 88.73%
Arena-Hard 45.36%
Auto-Arena 92.14%
As an example, we add a new participant (Llama-3-70B) to the existing 9-model ranking. It battles with $\lceil log_{2}(10)\rceil=4$ close opponents and Figure 5 shows how the Elo score changes throughout the rounds. Firstly, it is paired with Qwen-1.5 based on MMLU similarity and wins, which results in a very high Elo score, even above GPT-4. Then, it is paired with GPT-4, the closest opponent in Elo score. After losing, it is paired with the other opponents who are close in Elo scores, Command-R+ and Claude-3-Haiku. Eventually, the score stabilizes at second place. This process lets the new candidate battle with a reasonable fraction of close opponents and makes the final ranking stable without disrupting the other participants, whose score distribution remains similar before and after the addition.
Using this scalable addition approach, we build a comprehensive leaderboard by adding 6 new models to the existing tournament of 9 LLMs, resulting in a final ranking of 15 models. Figure 5 shows the overall Elo scores by Auto-Arena on the 15 models. Table 4 shows the Spearman correlations after expansion. Auto-Arena remains the method most aligned with human preferences by a margin of 3.41%, showing the state-of-the-art alignment of 92.14%. Therefore, Auto-Arena is generalizable and robust for maintaining a leaderboard for many LLMs.
4.2 Easy Extension to Other Domains and Languages
@@
As Auto-Arena of LLMs is fully automatic, it can be easily adapted to evaluate LLMs in other domains or languages. As case studies, we conduct a tournament in Chinese on models that are claimed to have multi-lingual proficiency. The only adaption effort is translating the prompts into the desired languages. Then, the generated questions and peer battles will be in the desired languages. It is also possible to adapt the framework to another task or domain, the only effort is to change the “domain” specification in the examiner’s prompts (shown in Appendix A).
Figure 6 shows the Elo scores derived by Auto-Arena for the Chinese tournament on 11 models. As Chinese evaluation benchmarks are limited, we compare with the Chinese-only leaderboard on Chatbot Arena, which constitutes 10.36% of all collected votes. We include 7 models best-performing and newest models from each major model family in the top 20 list on Chatbot Arena. The Auto-Arena recovers their Elo scores with a correlation of 92.86%, verifying the reliability of the extension. In addition, as Chatbot Arena doesn’t include proprietary Chinese LLMs, we add 4 popular Chinese LLMs, which are GLM [^3] https://open.bigmodel.cn/, SenseChat [^4] https://platform.sensenova.cn/home, Minimax [^5] https://platform.minimaxi.com/examination-center/text-experience-center, and Wenxin [^6] https://cloud.baidu.com/wenxin.html. We notice that the models claimed to have Chinese proficiency, such as Qwen-1.5, indeed score higher on this leaderboard compared to the English one.
5 Investigation of LLM’s Behaviors in Competitive Peer Battles
Peer Battles Make the Performance Gaps Become Visible
In the example shown in Figure 7, given a math question on infinite series, both candidate A (Claude-3-Haiku) and candidate B (GPT-4-Turbo) provide correct answers in the first round. However, as the debate deepens, the performance gap becomes more visible: Candidate B is able to provide a more elaborate and helpful response when explaining the theories behind the initial answer. In the ablation study without peer battles, the judges initially decided that it was a tie. However, after seeing the subsequent debates, they change to favoring assistant B. This example shows that the debate process indeed pushes the candidate LLM’s capabilities to the limit, testing deeper understandings and reasoning abilities. Moreover, as shown in the previous Table 2, the peer battles are indispensable for a robust and comprehensive evaluation.
@@ 5 Investigation of LLM’s Behaviors in Competitive Peer Battles
The example in Figure 9 shows excerpts of a peer battle around the question: “how many unique ways to arrange letters in ‘LETTER’.” Candidate A (powered by Yi-34B-Chat) gives a wrong initial answer as it miscounts occurrences for repeated letters and miscalculates factorials. The opponent B (powered by Claude-3-Haiku) quickly and precisely points out these two issues and skillfully raised a follow-up that targets A’s weaknesses: “how about the word ‘BANANA’?” Then, A still miscalculates factorials. We see that LLM candidates efficiently understand the rules of the competitive environment and can design targeted strategies to attack the opponent in order to win. In the peer battles, the debater agents display effective competition strategies, further probing the opponent’s weaknesses.
@@
6 Related Work
As LLMs evolve quickly, deriving trustworthy evaluations of their capabilities has become a challenge. Current evaluation methods can be divided into automatic evaluations and manual evaluations, such as Chatbot Arena (Chiang et al., 2024). We primarily focus on automatic evaluations as they deliver more timely feedback. Automatic evaluations mainly consist of static datasets with predefined metrics and model-based metrics. Static datasets with predefined metrics, such as MMLU (Hendrycks et al., 2021a), GPQA (Rein et al., 2023), and Open-LLM-Leaderboard (Beeching et al., 2023) consist of expert-annotated question-answer pairs. Then, the models are evaluated based on performance metrics such as accuracy. However, as they only evaluate closed-form answers, they are inflexible in evaluating open-ended responses. Moreover, the static datasets may eventually become exposed to the internet and could lead to contamination concerns (Ravaut et al., 2024).
On the contrary, static datasets with model-based metrics offer a flexible, low-cost and fast evaluation paradigm (Chang et al., 2024b). Studies have verified that LLMs can provide unbiased (Ning et al., 2024; Chu et al., 2024), high-quality (Lin & Chen, 2023) metrics comparable to human evaluations (Dubois et al., 2024a; Zheng et al., 2023). Among them, MT-Bench (Zheng et al., 2023) and AlpacaEval (Dubois et al., 2024a) use LLM-as-a-judge to ask GPT-4 to compare model responses to a static dataset of questions. The model’s judgments achieve over 80% agreement with human preferences, proving the usability of using LLMs to evaluate response quality. Language-Model-as-an-Examiner (Bai et al., 2024) asks an LM examiner to construct knowledge-intensive questions within its memory, interact with the candidate in a series of follow-up queries, and rate the responses on dimensions including accuracy and factuality. KIEval (Yu et al., 2024) also incorporates an LLM-powered “interactor” role to examine deep comprehension of knowledge, which is shown to mitigate contamination issues on static datasets. However, such single-judge evaluations require the examiner to interact with each candidate parallelly, creating computational overheads and limiting the scope of queries. They also suffer from single-model bias, including bias towards LLM-generated summaries (Liu et al., 2023), inflated scores in multilingual evaluation (Hada et al., 2023), verbosity bias (Dubois et al., 2024a), and difficulties when evaluating candidates with close performance (Shen et al., 2023). Therefore, there have been studies on employing multi-agent evaluation to mitigate single-model bias. For example, DRPE (Wu et al., 2023a) uses multi-roleplayer prompting to mimic different roles with the same LLM and integrate outputs as votes for the final results. PRD (Li et al., 2023a) allows two LLMs to discuss an evaluation and assigns higher voting weights to the LLM reviewers with stronger capabilities. They show that the multi-agent approach effectively mitigates single-model bias. This line of work is similar to our “LLM judge committee” component. However, they are still limited to static datasets and specific domains.
Outside the domain of LLM evaluations, some works study competitive behaviors in multi-agent LLM systems, which is relevant to the peer battles in Auto-Arena. LM vs LM (Cohen et al., 2023) shows that LLM cross-examinations can effectively discover factual errors. Debate (Du et al., 2023) shows that multi-agent debate can improve factuality and reasoning. In MAD (Liang et al., 2023), LLM-debate can encourage divergent thinking, which helps tasks that require deep levels of contemplation. Khan et al. (2024) shows that even non-expert weak LLMs can supervise expert LLMs if we allow the two LLM experts to engage in debates. Moreover, Zhao et al. (2023) and Gu et al. (2024) show interesting case studies where LLMs are engaged in simulated competitive environments and demonstrate human-like strategies.
@@ 7 Conclusions
In this paper, we innovatively design a completely automatic evaluation framework: Auto-Arena. By using LLM agents to generate questions, employing LLM candidates in peer battles, and evaluating responses using LLM committee discussions, Auto-Arena delivers timely and trustworthy evaluations and automates the evaluation process in an end-to-end way. In the extensive experiments, Auto-Arena achieves the highest correlation with human preferences, despite requiring zero human efforts. It is easily adaptable to other domains and resources, promoting the inclusiveness of AI system evaluations. The peer battles also demonstrate several interesting LLM behaviors in competitive environments, including attacking and learning from the opponents.
Appendix A Prompts Used
@@
A.1 Prompts to Examiner agent
@@
A.2 Prompts to Peer Battle Candidates
@@
This is the first prompt for the peer battle candidates. When possible, it is included as a system prompt. The action guide prompts are included in Table 6, where the actions are determined by the round and turn as illustrated in Figure 2.
@@ A.3 Prompts to Judges
This is the prompt for judges for discussion:
@@
Appendix B Example Questions Generated
@@ 2. What is the integral of ($x^{2}+2x+2)/(x^{3}+3x^{2}+3x+1)dx$?
Appendix C Contamination Analysis
Table 7: Average Contamination Percentages of Benchmarks.
Detection Method Ours MMLU ARC Challenge HellaSwag
GPT-4 Style (Substring Match) $\downarrow$ 2% 42% 33% 18%
Playtus Style (Sentence Similarity) $\downarrow$ 28% 41% 35% 43%
The design in the question-generation and peer-debate process ensures that contamination is minimized. Data contamination refers to the possibility of test instances showing up in pre-training or Supervised Fine-tuning data.
Question-generation:
As we generate the questions automatically, we reduce the risk of test instances being eventually exposed to the open web, which can happen in static datasets. Alleviation of data contamination is often shown to be an advantage of such dynamic and frequently updated evaluation frameworks (Li et al., 2023b).
Peer Debate:
Peer debate ensures that we evaluate the entire debate instead of simple question-answers, which further reduces contamination. During debates, the models are evaluated on comprehensive and deep abilities, such as planning the strategies, pointing out flaws of the opponents, and drafting further questions. Such interactive evaluation frameworks are shown to reduce contamination (Yu et al., 2024; Bai et al., 2024).
Besides the design choices, we conduct a contamination analysis to compare the contamination percentage of Auto-Arena debate questions and test questions in popular benchmarks. Specifically, we use two types of contamination detection metrics:
  1. 1.
    The string match metric as in GPT-4 (OpenAI et al., 2024), where a match is identified if any of three 50-character randomly sampled substrings from the evaluation data point (or the entire string if it is shorter than this) is a substring of the training set. If so, we mark the point as contaminated.
  2. 2.
    The sentence embedding similarity metric as in Platypus (Lee et al., 2024), where a question is deemed contaminated if it has a cosine similarity (using Sentence Transformer (Reimers & Gurevych, 2019) embeddings) greater than 80% against any training item. This detection method is more robust to rephrases, which ensures that we can detect cases where the LLMs are simply rephrasing existing questions on the web.
Although we do not have access to the training data, LLMs mostly use public web data for pre-training (Raffel et al., 2020; Brown et al., 2020; Touvron et al., 2023). Therefore, we approximate it with the Bing search API: If verbatim test examples appear online, it likely indicates inclusion or exposure to the training data. This procedure is also followed by Li et al. (2024) for detecting contamination.
The ablation is conducted as follows: Firstly, we randomly sample 100 questions from the testset. As baselines, we use 3 popular evaluation benchmarks: MMLU (Hendrycks et al., 2021a), ARC Challenge (Clark et al., 2018), and HellaSwag (Zellers et al., 2019). For each question, we get the top 10 search result snippets on the Bing search API. If the question is deemed as contaminated by the detection method (mentioned above) against any of the 10 snippets, it is marked as contaminated.
The percentages of contaminated test instances is reported in Table 7. We can observe that Auto-Arena, by generating fresh questions, does alleviate the contamination issue. Compared to static datasets, Auto-Arena’s contamination percentage (2%) according to the exact match is significantly lower. When using the sentence similarity metric, we can effectively detect whether generated questions are just rephrases of existing questions. The percentage is largely reduced by 7% to 15% compared to other benchmarks.
Appendix D Synthetic V.S. Real-Life Questions
In this section, we try to show the generalizability of the synthetic questions in Auto-Arena to real-life questions.
Design:
The generated questions resemble real-world queries by design. In the question generation prompt, we specifically ask the examiner to draft questions that are “likely for a user to ask in real life”. From Appendix B, we could also observe the similarity of the synthetic questions to real-life queries.
Table 8: Human Evaluation on Synthetic Questions and Real Questions.
Volunteer 1 Volunteer 2
Correct 27.1% 38.9%
Incorrect 27.1% 11.9%
Cannot Tell 45.8% 49.2%
Agreement -0.11
Human Study:
To show that the generated queries are similar to real-life ones, we conduct the following human study. We compare 30 synthetic questions by Auto-Arena and 30 real-life questions. A human user is asked to look at a question randomly drawn and decide whether he/she believes that it is AI-generated, Real-Life, or if he/she cannot tell. The questions are collected in the Math category, where the 30 real-life ones are taken from MT-Bench (10 questions, drafted by experts), AMC-8 (4 problems, from the 2024 math competition), and AGI-Eval (16 math questions collected from college entrance exams). Two volunteers who are frequent users of LLMs and are familiar with AIGC participated. We report their respective results and agreement in Table 8. We can observe that humans cannot tell if the problems are synthetic almost half of the time. The user accuracy (correct percentages) is also low. We calculate the Cohen’s Kappa agreement between the two users, which is -0.11. The agreement score shows that there is less agreement than random chance. The big divergence between human annotators’ responses also shows subjectivity and uncertainty in the judgments. Therefore, we conclude that humans most likely cannot tell whether questions are synthetic or real-world, indicating small differences.
Table 9: Ablation Results on Synthetic Questions and Real Questions.
Questions GPT-4 Win Rate Claude-3 Win Rate
Synthetic Questions 80.00% 20.00%
Real-life Questions 75.86% 24.14%
Ablation Study:
To validate the results’ generalizability with real-world datasets, we conduct an ablation study comparing Auto-Arena’s evaluation performances on real-life questions and synthetic questions. Specifically, we asked 2 candidates (GPT-4-Turbo-0409 and Claude-3-Haiku) to debate around 30 synthetic math questions and 30 real-world math questions (collected as in the human study shown in Table 8). If the results are generalizable, we would observe that the win rates of each model should be similar. The results are shown in Table 9. From the results, we can observe that the win rates of each model only differ by 4% on synthetic and real datasets, which shows consistent evaluation performances, validating the use of synthetic problems.
Aside from the supporting studies, the use of synthetic questions for evaluation has also been established as common practice. The Mathematics dataset (Hendrycks et al., 2021b) already uses synthetically generated math questions, where they note many advantages, such as the ease of providing a larger number of examples, the precise controls over difficulty levels, and the ease of testing generalization (since one can precisely vary different axes of difficulty in different question types). LMExamQA (Bai et al., 2024) also uses an LLM to generate questions in different domains. KI-Eval (Yu et al., 2024) asks an LM-powered interactor to generate questions. The list goes on. Using synthetic questions has become the common norm in NLP evaluation. Moreover, extensive experiments in Auto-Arena show high correlations with human results, which also demonstrates the alignment with real-world usage.
Appendix E Ablation study on Self-Enhancement Bias of the Question Generation Stage
Table 10: Ablation Results on Self-Enhancement Bias for Question Generator.
Questions GPT-4 win rate Haiku win rate
GPT-4 Generated Questions 80.00% 20.00%
Haiku Generated Questions 76.92% 23.08%
We attempt to reduce self-enhancement bias of the question generation stage with explicit designs: Firstly, during question generation, we do not disclose to the examiner that it will participate in this tournament and we do not ask the examiner to generate only questions that can be solved by itself. Secondly, the peer-debate process further reduces bias in initial question generation: Debating ensures that candidates are evaluated not only on their response to the initial question, but also in more comprehensive and deeper abilities, such as strategizing, criticizing the opponent, and drafting questions. In other words, answering the initial question well does not necessarily win a whole debate. In the debate design in Figure 2, candidates also have a ”raise“ action, where they ask questions to the opponent. This process essentially decentralizes the question-generation process.
To systematically examine whether self-enhancement bias is present. We conduct an ablation study: We examine enhancement bias with 2 models as an example: GPT-4 (GPT-4-turbo) and Haiku (Claude-3-Haiku). Firstly, we ask GPT-4 and Haiku to generate 30 math questions separately. Then, we conduct peer debates between the two candidates (GPT-4 and Haiku) on both sets of questions and evaluate results with the best-5-LLM committee as in the main experiments.
We evaluate the performance differences from the evaluation results: If self-enhancement bias is low, the ranking achieved should remain the same. In other words, the weaker model will always lose, even on the questions generated by itself.
The ablation results are shown in Table 10. From the results, we can observe that, in both sets of generated questions, the GPT-4 win rate remains significantly higher than the Claude-3-Haiku win rate. Even if some limited extent of self-enhancement bias is present, the result difference is significant enough to reach the correct ranking.
Appendix F Inter-judge Agreement